OpenStudy (anonymous):

how do i find the geometric mean between 4 and 1/4

OpenStudy (anonymous):

lol

OpenStudy (rsadhvika):

multiply them

OpenStudy (rsadhvika):

then take square root

OpenStudy (anonymous):

k

hartnn (hartnn):

here a =4, c=1/4

OpenStudy (anonymous):

would i set it up the same way or?

hartnn (hartnn):

exactly same way

OpenStudy (anonymous):

lik 4/b =1/4 ??

hartnn (hartnn):

nopes, its was a/b= b/c so, 4/b = b/ (1/4)

OpenStudy (anonymous):

ok ok, hmmm one sec lol

OpenStudy (anonymous):

its 1 ?

hartnn (hartnn):

yup! :)

OpenStudy (anonymous):

i feel smart ^~^ never gotted these right :) b4 xD

OpenStudy (anonymous):

thankyou so much :))

hartnn (hartnn):

welcome ^_^ you'll always get these right from now ;)

OpenStudy (anonymous):

\(^~^)/